Consider the utility function u (c, o) = oc (o = leisure; c = consumption), determine the optimal amount of consumption and leisure if the consumer can work at most 24 hours, the hourly wage is 5 and the price of each unit of consumption is 2. There is no initial income endowment.
Answer #1

Answer Lets first form budget constraint 0+ w = T o = Leisure where and T = w= hours worked Time available =24 > or w=T - W=